# LLM-Assisted Static Analysis for Detecting Security Vulnerabilities

### v1

@@ 6 Related Work
LLMs for software engineering. Researchers are increasingly combining LLMs with software engineering tools for challenging tasks such as fuzzing [26, 45], program repair [44, 22, 43], and fault localization [46]. While we follow a similar direction, to the best of our knowledge, our work is the first to combine LLMs with static analysis to detect security vulnerabilities via whole-project analysis. Recently, LLM-based agents such as AutoCodeRover [47] and SWE-Agent [42] are also pushing the boundaries on whole-project repair. Hence, in the future, we plan to explore a richer combination of tools in IRIS to further improve the performance of vulnerability detection.
@@ 7 Conclusion and Limitations
We presented IRIS, a novel neuro-symbolic approach that combines LLMs with static analysis for vulnerability detection. We also curated a dataset, CWE-Bench-Java, containing 120 security vulnerabilities across four classes in real-world projects. Our experimental results show that IRIS can significantly surpass state-of-the-art static analysis tools, like CodeQL, in detecting critical vulnerabilities, even when using smaller LLMs. Further, IRIS’s context filtering techniques also drastically reduce the number of false positives. Overall, our results show that effectively combining LLMs with static analysis leads to more effective analysis and reduces the developer burden. However, there are still many vulnerabilities that can yet be detected by this approach. Hence, future approaches may explore a tighter integration of these two tools to further improve performance.
Limitations. Determining whether a vulnerability is detected is a difficult task. Hence, we rely on manual validation to find vulnerable method locations in each case, limiting the size of CWE-Bench-Java. However, because each project is still quite large, we believe CWE-Bench-Java will remain a challenging benchmark for future works. IRIS makes numerous calls to LLMs for specification inference and filtering false positives, increasing the potential cost of analysis. However, as our results show, even smaller models, like Llama 3 8B and DeepSeekCoder 7B, can perform well on these tasks and can be potentially deployed locally for a project. For any given project, the cost of labeling will consolidate quickly across versions, making IRIS a viable choice for potential integration into CI/CD pipelines. While our results on Java benchmarks are promising, it is unknown if IRIS will perform well on other languages. We plan to explore this further in future work.
@@ A.4 Visualization of Metrics
We provide a visualization of our VulDetected metric in Fig. 7. For evaluation, we assume that the label for a project is provided as a set of crucial program points where the vulnerable paths should pass through. In practice, these are typically the patched methods that can be collected from each vulnerability report. As illustrated in Fig. 7, if at least one detected vulnerable path passes through a fixed location for the given vulnerability, then we consider the vulnerability detected. Suppose is the set of detected and/or filtered paths from prior stages, the metrics, including #Vuls the total number of detected vulnerabilities across a set of projects, is formally defined as:
A.5 Discussion of Precision Metrics
In contrast to VulDetected, computing the precision of the results is more challenging. For instance, even if a detected code path does not intersect with a fixed file or method, it may actually point to a true vulnerability (e.g., a different CVE in the same version) in the project. Moreover, even if there is a Hence, manual analysis is required to compute precision.
@@ C.1 Experimental Settings
We use CodeQL version 2.15.3 as the backbone of our static analysis.
@@ C.4 Details of Selected LLMs
Gemma 7B | google/gemma-1.1-7b-it |
Mistral 7B | mistralai/Mistral-7B-Instruct-v0.2 |
C.5 Ablation Studies
Table 7 presents additional results when using either only the source or sink specification from an LLM in IRIS. For this experiment, we only use the results with GPT 4 for comparison. Each row present the average number of paths and number of detected vulnerabilities per CWE, for each combination. We observe that omitting either source or sink specifications inferred by GPT 4 causes a drastic reduction in overall recall. For CWEs 22, and 78, inferring additional source specifications seems to be more beneficial for improving recall (row 2). This is because for CWE 22 (path traversal), the sinks are typically File related methods (e.g., FileOutputStream) and for CWE 78, sinks are usually system calls (e.g., Runtime.exec) – many of which are available in CodeQL’s specifications. Howver, the potential sources for these CWEs can be more diverse and hence missing in CodeQL. On the other hand, for CWE 79 (Cross-Site Scripting) and CWE 94 (Code Injection), the sinks can vary widely including XML/HTML manipulation APIs or deserialization-related APIs, many of which are missing in existing specifications. Hence, sink specifications are more useful for these two CWEs.
Metrics | #Detected Vuls | ||||||||
CWE | 22 | 78 | 79 | 94 | |||||
Combination | #Paths | #Vuls | #Paths | #Vuls | #Paths | #Vuls | #Paths | #Vuls | Total |
172 | 22 | 3 | 1 | 31 | 4 | 1 | 0 | 27 () | |
202 | 28 | 7 | 3 | 43 | 5 | 0 | 0 | 36 () | |
82 | 10 | 57 | 1 | 335 | 9 | 369 | 4 | 24 () | |
760 | 34 | 529 | 7 | 2139 | 17 | 1590 | 11 | 69 |
@@ C.6 Statistics of Unique and Recurring Specifications
C.7 Analysis Runtime
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ 6 Related Work
LLM-based approaches for software engineering. Researchers are increasingly combining LLMs with program reasoning tools for challenging tasks such as fuzzing [27, 46], program repair [45, 22, 44], and fault localization [47]. While we are on a similar direction as [28], to our knowledge, our work is among the first to combine LLMs with static analysis to detect application level security vulnerabilities via whole-project analysis. Recently, LLM-based agents such as AutoCodeRover [48] and SWE-Agent [43] are also pushing the boundaries on whole-project repair. Hence, in future, we plan to explore a richer combination of tools in IRIS to further improve the performance of vulnerability detection.
@@ 7 Conclusion and Limitations
We presented IRIS, a novel neuro-symbolic approach that combines LLMs with static analysis for vulnerability detection. We curate a dataset, CWE-Bench-Java, containing 120 security vulnerabilities across four classes in real-world projects. Our results show that systematically combining LLMs with static analysis significantly improves upon traditional static analysis alone in terms of both detected bugs and the alleviation of developer burden.
Limitations. There are still many vulnerabilities that IRIS cannot detect. Future approaches may explore a tighter integration of these two tools to improve performance. In addition, IRIS makes numerous calls to LLMs for specification inference and filtering false positives, increasing the potential cost of analysis. While our results on Java are promising, it is unknown if IRIS will perform well on other languages. Moreover, there is still a gap between the IRIS generated report and the report that the developers would like to see. We plan to explore this further in future work.
@@ A.4 Visualization of Metrics
We provide a visualization of our VulDetected metric in Fig. 10. For evaluation, we assume that the label for a project is provided as a set of crucial program points where the vulnerable paths should pass through. In practice, these are typically the patched methods that can be collected from each vulnerability report. As illustrated in Fig. 10, if at least one detected vulnerable path passes through a fixed location for the given vulnerability, then we consider the vulnerability detected. Let be the set of detected paths for each project from prior stages. The vulnerable paths inside project is given by:
@@ C.1 Experimental Settings
We use CodeQL version 2.15.3 as the backbone of our static analysis. We have patched CodeQL with an additional feature that augments the Dataflow edge between throw statement and its closest surrounding try-catch block. We use this CodeQL pull request as the base of our patch.
@@ C.4 Details of Selected LLMs
C.5 Previously Unknown Bugs Found by IRIS
We have applied IRIS with GPT-4 to 30 open-source Java projects in order to find previously unknown vulnerabilities. The detailed statistics, at the time of this paper’s submission, is listed in Table 7. We inspect vulnerabilities in detail, and will only file when we think that there is a vulnerability with high severity and an easily accessible attack surface. 3 out of the 4 filed vulnerabilities are filed on HackerOne to abide to the rules of the individual open-source project, which prohibit us to disclose the details before the vulnerability is fixed. The remaining one is found in the Java project alluxio 2.9.4, which we disclose and discuss in detail in Fig. 8. For the already filed vulnerabilities, we are committed to work till the end to make sure that they are fixed in the open-source project. The paper will be updated when they are fixed by the developer.
#Projects | 30 |
#Projects where IRIS reports alarm | 16 |
Average #alarms per investigated project | 5.78 |
#Projects with potential vul | 4 |
#Vuls with potential attack surface | 6 |
#Vuls filed for developer confirmation | 4 |
#Vuls of CWE-22 | 3 |
#Vuls of CWE-78 | 0 |
#Vuls of CWE-79 | 0 |
#Vuls of CWE-94 | 1 |
#Developer Response | 0 |
#Developer Confirmation | 0 |
#Fixed Vulnerability | 0 |
Our experience with finding unknown vulnerabilities using IRIS is pleasant, especially given that we have built a custom web-based interface for us to inspect and triage the vulnerabilities. A screenshot of the user-interface is illustrated in Fig. 11. We may quickly scan over the vulnerabilities by looking at their source, sink, and LLM explanation of whether it is a bug. Additionally, all the dataflows and nodes have hyperlinks to Github and the source code, making it easy to navigate through the codebase. Overall, it does not take too much time to dive deep into each vulnerability. Anecdotally, the first two real CWE-22 vulnerabilities are found among the first two projects that we inspected.
@@ C.6 Statistics of Unique and Recurring Specifications
C.7 Statistics of Inferred Taint Specifications
We show the statistics of inferred taint specifications in Table 10. As shown by the percentage, GPT-4 generates smaller set of sources and sinks than smaller-scale LLMs like DeepSeekCoder 7B.
CWE | #Cand. | GPT-4 | DSC 7B | ||
%S | %N | %S | %N | ||
22 | 130,974 | 2.03% | 1.90% | 4.27% | 4.01% |
78 | 25,605 | 4.73% | 3.37% | 3.67% | 3.33% |
79 | 37,138 | 5.69% | 4.69% | 4.28% | 4.56% |
94 | 36,325 | 5.12% | 7.83% | 6.11% | 6.21% |
Total | 230,042 | 3.41% | 3.45% | 4.50% | 4.37% |
Appendix D Analysis Runtime
@@
22 | uima-uimaj | 226.81K | 2m | 5.66K | 103 | 176 |
22 | keycloak | 614.82K | 12m | 13.34K | 325 | 252 |
22 | glassfish | 1.19M | 5m | 12.19K | 293 | 346 |
22 | graylog2-server | 382K | 4m | 13.3K | 227 | 171 |
22 | mina-sshd | 130.14K | 1m | 3.64K | 52 | 120 |
@@
79 | DSpace | 237.33K | 2m | 3.67K | 347 | 320 |
79 | keycloak | 615.6K | 3h | 13.37K | 606 | 461 |
79 | keycloak | 615.6K | 3h | 13.37K | 606 | 461 |
79 | xwiki-commons | 105.92K | 1m | 3.94K | 244 | 151 |
79 | xwiki-commons | 105.94K | 1m | 3.94K | 244 | 151 |
@@
79 | xwiki-commons | 106.87K | 1m | 3.99K | 254 | 161 |
79 | jspwiki | 158.7K | 1m | 2.22K | 176 | 126 |
79 | keycloak | 617.15K | 4h | 14.04K | 643 | 479 |
79 | xwiki-commons | 107.09K | 1m | 3.03K | 209 | 143 |
@@
94 | dubbo | 175.63K | 2m | 6.73K | 342 | 383 |
94 | incubator-dubbo | 96.35K | 1m | 3.68K | 194 | 255 |
94 | spring-security | 57.34K | 1m | 2.43K | 192 | 234 |
94 | kubernetes-clie | 806.35K | 3m | 2.33K | 93 | 130 |
@@
94 | nifi | 993.76K | 25m | 57 | 2 | 11 |
94 | rocketmq | 108.39K | 2m | 3.4K | 117 | 164 |
94 | nifi | 1.01M | 27m | 261 | 27 | 24 |
94 | rocketmq | 197.78K | 2m | 6.28K | 205 | 252 |
94 | dolphinschedule | 154.95K | 4m | 5.78K | 229 | 353 |
94 | dolphinschedule | 154.95K | 4m | 5.78K | 229 | 353 |